Circular No. 6909

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998cd IN ANONYMOUS GALAXY J. Mueller reports her discovery of an apparent supernova (mag ∼ 17.5)

on a plate taken on May 17 with the 1.2-m Oschin Schmidt Telescope in the course of the second Palomar Sky Survey. R. Gal obtained a CCD frame on May 18 with the Palomar 1.5-m telescope that confirms the presence of the object. G. V. Williams, Smithsonian Astrophysical Observatory, provides the following reduction of positional measures by Mueller for SN 1998cd: α = 10^{h}45^{m}04^{s}.16, δ = +39^{o}25′12′′.5 (equinox 2000.0); this corresponds to an offset 8′′.2 east and 5′′.8 south from the galaxy’s center. The object does not appear on original Sky Survey prints or on a IIIa-J Sky Survey plate taken on 1991 Feb. 9 by C. Brewer and J. D. Mendenhall.

GRB 980515 M. Feroci and L. Piro, Istituto Astrofisica Spaziale, CNR, Rome; M. R.

Daniele, G. Gennaro, S. Rebecchi, and G. Celidonio, BeppoSAX Science Operation Center, Rome; L. A. Antonelli, BeppoSAX Science Data Center, Rome; and J. Heise, Space Research Organization of the Netherlands, Utrecht, report: “The BeppoSAX Wide Field Camera error box of GRB 980515 was observed with the BeppoSAX Narrow Field Instruments (NFI) starting on May 16.1178 UT (10 hr after the burst). A previously unknown x-ray source, 1SAX J2116.8–6712, was detected by the MECS instruments (units 2 and 3) at α = 21^{h}16^{m}49^{s}, δ = −67^{o}11′.8 (equinox 2000.0; preliminary error radius 1′.5 due to the poor statistical quality). This position is 4′.9 from the centroid of the WFC position (GCN 73). The source 1SAX J2116.8–6712 exhibits a moderate variability over the two-day NFI observation. There is no clear evidence for a monotonic decaying behavior. The average count rate was 0.0026 ± 0.0006 count/s (2-10 keV) in the two MECS units, corresponding to (1.6 ± 0.4) × 10^{−13} erg cm^{−2} s^{−1} (assuming a Crab-like energy spectrum). At this time, 1SAX J2116.8–6712 is the only candidate for the x-ray afterglow of GRB 980515.”

SUPERNOVA 1998aq IN NGC 3982 CCD magnitude estimates, unfiltered unless otherwise noted: May

7.862 UT, 12.9 (G. Masi, Ceccano, Italy); 8.915, V = 12.53 (D. Hanžl, Brno, Czech Republic); 8.99, 12.8 (L. Boschini, Monte di Nese, Bergamo, Italy); 9.898, 13.0 (Masi); 16.899, 13.3 (Masi).

1998 May 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green